OpenStudy (anonymous):

how do i solve |2/3x+5|=1/4

OpenStudy (anonymous):

im' doing it wrong

terenzreignz (terenzreignz):

This. \[\huge \left|\frac{2}{3x+5}\right|=\frac14\]?

OpenStudy (anonymous):